OpenStudy (anonymous):

Solve for x: (x-16)/3 = 3x A:x = −7 B:x = −13 C:x = 8 D:x = 2 I'm confused with the two x's in the problem and if it's a little confusing the / stands for the fraction sign not the division sign.

OpenStudy (anonymous):

Multiply both sides by three to get rid of the fraction

OpenStudy (anonymous):

Then you have :x-16=9x. Subtract an x from both sides giving you: -16=8x. Finally divide both sides by 8 resulting in : -2=x

OpenStudy (anonymous):

(x-16)/3=3x 9x=x-16 9x-x=-16 x=-2

OpenStudy (anonymous):

thankkk you for the help I appriciate itt :D